Please help me with this questions!
Consider a function f(x) = x2+x+1
1) Specify the integrated function F(x) for the given f(x).
2) Find the value of the area between 1 and 2 under the integrated function F(x)

Answer by stanbon(75887)

Consider a function f(x) = x2+x+1
1) Specify the integrated function F(x) for the given f(x).
F(x) = (1/3)x^3 + (1/2)x^2 + x
----------------------------------
2) Find the value of the area between 1 and 2 under the integrated function F(x)
F(2)-F(1) = [(8/3)+(2)+2]-[(1/3)+(1/2)+1] = [(7/3)+(3/2)+1] = 29/6 = 4.83333..
=====================================
Cheers,
Stan H.